OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (amistre64):

how far have you gotten on it?

OpenStudy (amistre64):

this is the help, come back and lets finish this up

OpenStudy (anonymous):

ok...

OpenStudy (amistre64):

the first thing we want to do is get rid of those paranthesis. Do you know how we can do that?

OpenStudy (anonymous):

subtracting a number from both sides?

OpenStudy (amistre64):

thats going to play a part in the solution later, but for now, let try multiplying the number on the outside to everything inside the paranthesis. whats 4 times x equal? whats 4 times -6 equal?

OpenStudy (amistre64):

and on the other side of the = sign: -2(x-3) whats -2 times x equal? whats -2 times -3 equal?

jhonyy9 (jhonyy9):

4x-24=-2x+6 6x=30 x=30/6 x=5

jhonyy9 (jhonyy9):

hope that is understandably

OpenStudy (anonymous):

4 times -6 equals is -24. -2 times -3 equals 6

OpenStudy (amistre64):

thats right so far; the ones with the xs just get squished together. 4 times x = 4x -2 times x = -2x so lets rewrite it out again to see what it looks like: 4(x-6) = -2(x-3) 4x -24 = -2x +6 hopefullt this is looking more doabl in this form

OpenStudy (amistre64):

now we can do adding and subtracting from side to side :)

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

4x -24 = -2x +6 how would we get rid of the -24? how would we get rid of that -2x?

OpenStudy (amistre64):

in other wirds, we want to put all the xs to one side and all the not xs to the other

OpenStudy (anonymous):

subtract the 4x from both sides.the 4x would equal out and then 24-4x

OpenStudy (amistre64):

youve got the right idea, a bit more practice would make it more solid tho. lets see what subtracting 4x will do for us tho. 4x -24 = -2x +6 -4x -4x ----------------- 0 -24 = -6x + 6 so; -24 = -6x + 6 is something we can work with. but i think we need to get rid of the +6 as well

OpenStudy (amistre64):

how do we get rid of the +6 ?

OpenStudy (anonymous):

add it to one of the sides?

OpenStudy (amistre64):

lets try subtracting it from both sides; just to keep it all balanced. -24 = -6x + 6 - 6 -6 -------------- -30 = -6x

OpenStudy (amistre64):

can you see the solution yet? or do we need to do one more step?

OpenStudy (anonymous):

one more step.

OpenStudy (amistre64):

:) that -6 sticking to the x is the bad guy that we need to get rid of; but there is no more adding or subtracting that we can do; to undo the multiplication we have to divide it off. does that make sense?

OpenStudy (amistre64):

in other words; -6x/-6 = x would you agree?

OpenStudy (anonymous):

Yeah, i think so.

OpenStudy (amistre64):

lets try it out and see :) -30 = -6x /-6 /-6 ----------- -30/-6 = x and there we have it: x = -30/-6 can you simplify that any more?

OpenStudy (anonymous):

no, we cannot.

OpenStudy (amistre64):

well, lets look at it a little bit more just to make sure: -30/-6 , hmm 2 negatives makes a positive 30/6 now i know this looks like a fraction but all a fraction is is dividing so we can read this as: what is 30 divided by 6 equal?

OpenStudy (amistre64):

if dividing isnt a strong point yet, we can go back a step: -6x = -30 this reads as: -6 times (what?) equals -30.